OpenStudy (anonymous):

The length of a rectangle is six times its width. If the perimeter of the rectangle is 98ft , find its area.

OpenStudy (anonymous):

2l+2w=98 2(l+w)=98 l+w=49 6w+w=49 7w=49 w=7 the area is (w)(6w)=294 ft